OpenStudy (anonymous):

Figure out the pattern below and then predict the next term. 89, 77, 65, 53, 41, ... 29 30 31 32

OpenStudy (michele_laino):

hint: we can write this: \(77=89-12\) \(65=77-12\) \(53=65-12\) \(41=53-12\) every time I subtract \(12\) units, so what is: \(41-12=...?\)